# Estimation of motion blur kernel parameters using regression convolutional neural networks

### v2

@@ 3.3 COCO Blurred Dataset
(a) Number of unique COCO images per angle in training and testing/validation datasets.
(b) Number of unique COCO images per length in training and testing/validation datasets.
Figure 4: Representation of COCO images versus length and angle.
@@
We use the 2014 COCO dataset [23] to create a blurred dataset for training and validating a model for blind estimation of length and angle given only a blurry image. The 2014 COCO dataset consists of a training dataset with 82,783 images and a validation set of 40,504 images. We use the training dataset to create a blurred training dataset and we split (with no overlap) the validation dataset to create blurred validation and test datasets.
@@
We create the blurred dataset by creating a blur generator that loops through the labeled pairs $(r_{u},\phi_{u})$, creating the corresponding blur kernel, and convolving the blur kernel with an image from the COCO dataset. We choose a random COCO image (without replacement) for each length $r_{u}$ to provide a variety of images in the blurred dataset. We encounter a dataset imbalance in lengths and/or angles represented in the blurred dataset due to the uneven representation of angles for shorter lengths (see Figure 3(b)). We improve the imbalance by creating a minimum of 175 blurred images per length. This implies that we end up using more unique COCO images for shorter blur lengths. Each run of the blur generator creates 21,789 blurred images. The creation of the training set used 12 parallel threads of the blur generator, each operating on a subset of the COCO dataset, creating a total of 261,468 blurred images. The validation and testing sets were each created with one run of the blur generator, creating a total of 21,789 blurred images for each. Figure 4 shows the number of unique COCO images versus length and angle for the training, validation, and testing datasets. We notice a spike in number of unique images for length $r=1$ and for angle $\phi=0$. This is due to the non-blurred images that are part of the dataset. We notice more unique images used for shorter lengths as the blur generator needs to loop through more images to complete the minimum threshold of images per length. We notice these small lengths create peaks in the angles of $\phi=[-45,0,45,90]$.
(a) Number of blurred images per angle.
(b) Number of blurred images per length.
Figure 5: Distribution of blurred images for angle and length.
@@
We note that there is an entanglement between angle and length, resulting in an inability to completely balance both length and angle in the blurred dataset. Figure 5 shows the distribution of labeled blurred images with respect to angle and length. Our choice to generate a minimum of 175 blurred images per length in the blur generator favors a more uniform length distribution [Figure 5(a)] This, however, results in peaks in the angle distribution [Figure 5(b)] due to the over-representation of certain angles for smaller lengths (see Figure 3). A choice to favor a more uniform angle distribution, however, would result in missing many shorter blur lengths, creating a more problematic data imbalance. While there are other choices that could be made regarding data imbalance here, we find that this blurred dataset can be used to train a network to accurately predict both length and angle.
@@ 4 Uniform Blur Prediction
We formulate blur length and angle estimation as a deep learning regression problem using the VGG16 [3] architecture as the backbone of our model. We trained VGG16 from scratch with TensorFlow’s default layer initializations to predict both the length and angle parameters for a linearly uniform motion blurred image. We modified the last layer of VGG16 to have two output nodes (corresponding to prediction of length and angle) with sigmoid activations. Since the sigmoid activation outputs values in the range $[0,1]$, we normalize length and angle to be in the range $[0,1]$ to train the model. The predicted length and angle parameters are re-scaled to their native ranges ($r\in[1,100]$, $\phi\in(-90,90]$) for validation.
Due to the fixed input size of VGG16 $(224\times 224)$, we created a TensorFlow data generator to randomly crop the blurred COCO images. We crop instead of resizing the image to maintain accuracy in the blur angle labels since a resize could change the aspect ratio of the image and thus the blur angle. If a blurred image was smaller than the input size for VGG16 then it was skipped and not used. In total there were 892 training, 5 validation, and 13 test images that were skipped. This results in less than $0.3\%$ of images skipped for each dataset.
We use the Adam optimizer to with a learning rate of 0.1 and epsilon of 0.1. We use the mean squared error (MSE) as the loss function. We empirically determined that a batch size of 50 minimized convergence issues. We train for 50 epochs, saving the weights for the best model throughout training and terminating training if the MSE performance has not improved within the previous 5 epochs. Training takes about 25 minutes per epoch and about 12 hours to fully train on an NVIDIA RTX-3090.
We performed multiple trainings of the network by adding different noise levels to each training. This resulted into four different models. The first model is trained with no noise and the other three models are trained with different levels of additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$. We added the noise in our data generator after the image was cropped and normalized to the range $[0,1]$. In testing we test each model with the noiseless test set and by adding the three noise levels to the test set for an additional three test runs.
@@ 5 Experiments and Results
In this section, we present results using the $R^{2}$ metric for evaluating the regression model. We present results of the effects of additive noise on our model’s predictions. Finally we compare with other methods [5, 11] and evaluate using the error ratio score as introduced by [24].
@@ 5.1.1 Accuracy of Parameter Estimation
In testing, we measure performance using the coefficient of determination, $R^{2}$. The coefficient of determination, $R^{2}$, measures the goodness of fit between actual known values of variable $y_{i}$ and estimated values $x_{i}$:
@@
where $\bar{y}$ is the mean of the known variables $y_{i}$ and $n$ is the number of samples [25]. The numerator $\sum_{i=1}^{n}(y_{i}-x_{i})^{2}$ is the sum of squares of the residual prediction errors and the denominator $\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ is proportional to the variance of the known data. A perfect model will have zero residual errors and thus an $R^{2}=1$. A naïve model that always predicts the average of the data $\bar{y}$ will have equal numerator and denominator and thus an $R^{2}=0$. Models that have predictions worse than the naïve model will have $R^{2}<0$.
@@ 5.1.2 Quality of Deblurred Image
In addition to studying the accuracy of blur kernel estimation, we measure the quality of the image deblurred using the estimated blur kernel in a non-blind deblurring method (see Section 5.3). To measure the quality of the deblurred image, we use the error ratio as presented in [24]. The error ratio is motivated by the fact that, even with a perfectly estimated blur kernel, one may not be able to perfectly predict the latent sharp image. The error ratio $E_{\hat{k}}/E_{k}$ is computed by considering the error between the true sharp image and the latent sharp image recovered using the estimated blur kernel $E_{\hat{k}}$ and the error between the true sharp image and the latent sharp image recovered using the true blur kernel $E_{k}$. An error ratio of 1 indicates that the images deblurred using the estimated and true kernel are identical.
Any error metric can be used to define the error ratio, but we use the sum of squared differences (SSD) as in [24]. The error ratios are generally presented as a cumulative histogram for error ratios binned in the range $[1,4]$. As noted in [24], SSD error ratios above two tend to indicate significant perceptually noticeable distortion present in the image deconvolved with the estimated kernel.
@@ 5.2.1 Noise-Free Predictions
(a) Scatterplot of estimated versus actual length, $R^{2}=0.9869$.
(b) Scatterplot of estimated versus actual angle, $R^{2}=0.9935$.
Figure 6: Scatterplots of estimated versus actual values for (a) length and (b) angle along with corresponding $R^{2}$ score. The individual estimates are represented as blue points and the best fit linear line is plotted in red.
A VGG-16 based model (see Section 4) was first trained on blurred images without any additive noise. Scatterplots of estimated versus actual length and angle are illustrated in Figure 6 along with the corresponding $R^{2}$ scores. We find the model to be highly accurate for prediction of both length and angle as demonstrated by the $R^{2}$ scores of 0.9869 and 0.9935, respectively. In Figure 6(a) we note a larger spread in estimated values as the length increases. This is not surprising, as larger blur lengths are expected to be more difficult to accurately estimate. In Figure 6(b) we note certain angles have a larger spread in estimated values. This is most notable for $\phi\in\{0,\pm 45,90\}$, but can be noted for other angles. This is due to errors in prediction for the smaller length kernels which have a limited set of angles. It is important to recall, however, that many of these incorrect angle predictions will result in a correct blur kernel. As an example, an image blurred with kernel parameters $(r,\phi)=(2,0)$ may have an angle prediction of $\hat{\phi}=27$. However, since blur kernels of length $r=2$ can only be represented by $\phi\in\{0,\pm 45,90\}$, the blur kernel created with parameters $(r,\phi)=(2,27)$ will generate the closest valid line of $(r_{u},\phi_{u})=(2,0)$.
@@ 5.2.2 Predictions with Additive Noise
We used the model from Section 5.2.1, trained on noise-free blurred images, and tested it on images with additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$, corresponding to signal-to-noise (SNR) values of $\{30,20,10\}$ dB, respectively. Results for this experiment are shown in the first row of Table 1 and Table 2 for length and angle prediction, respectively. We note the estimation of the length parameter is more susceptible to noise than angle, resulting in a complete failure of prediction ($R^{2}$ scores less than 0) for even the smallest level of additive noise, $\sigma^{2}=0.001$. We hypothesize that additive noise can alter the intensity distribution along the blur path in a blurry image, creating the appearance of artificially shorter or longer blur paths. Those blur paths, however, will likely retain more characteristics of their angle for the same level of noise.
We trained three additional models using noisy blurred images with additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$ and tested each of those models on noise-free and noisy images, i.e., for $\sigma^{2}\in\{0,0.001,0.01,0.1\}$. Results for those three models are in the second through fourth rows of Table 1 and Table 2 for length and angle prediction, respectively. We again note a higher sensitivity to noise in the prediction of length (Table 1) than angle (Table 2). We further note that the $R^{2}$ score decreases by $\sim 0.1$ for the model trained on the highest level of noise $\sigma^{2}=0.1$ compared to the noise-free model, but that the same model is robust to varying levels of noise. Finally we note that the models trained on noisy data appear to be robust to noise levels less than or equal to the noise level on which they are trained. This implies that a single model trained on a single noise level can yield accurate predictions even for smaller noise levels not seen in training.
In comparison to other methods that test under additive noise, most, e.g. [11, 7], test up to a Gaussian noise of $\sigma^{2}=0.01$ to simulate sensor noise. Krishnan et al. [5] tests noise of up to $\sigma^{2}=0.02$. Our model is tested up to $\sigma^{2}=0.1$ and demonstrates a higher tolerance for noise which can be an advantage when modeling atmospheric turbulence.
@@ 5.3.1 Deconvolution and Comparison Methods
We quantify the performance of our blur parameter estimation by using the expected path log likelihood (EPLL) method [26] to deconvolve the images using our predicted motion blur kernel and the ground truth blur kernel. We used the python implementation available at https://github.com/friedmanroy/torchEPLL. We replaced their Gaussian kernel with our linear motion blur kernel.
EPLL only accepts square, odd-sized blur kernels. This means we must zero pad our predicted parameter kernel to a square, odd-sized blur kernel. We zero padded the shortest side of the kernel on both sides to keep the kernel centered and padded to match the size of the longest side. If the longest side is even, however, this would result into an even-sized square kernel. We used a similar approach to [27] where we create four odd-sized kernels each zero padded with the line asymmetrically offset toward a different corner. We deconvolve the blurred image with each of the four kernels and average the four deconvolved images as the resulting deblurred image.
Additionally, we compare performance with other blur kernel estimation methods from Levin et al. [11] and Krishnan et al. [5]. Both Levin et al. [11] and Krishnan et al. [5] include methods for estimating the deblurred image. For fairness of comparisons, we use only the blur kernel estimate from both methods and use their blur kernel estimates in the same EPLL [26] deblurring framework as we use for our estimated blur kernels. For the method in [11], we used the Matlab implementation provided at https://webee.technion.ac.il/people/anat.levin/. We used the deconv_diage_filt_sps function and set a kernel prediction size of $(101,101)$. This is to allow prediction of the largest kernel size in our blurred dataset. It should be noted that the method in [11] estimates a blur kernel for each of the three channels in a color image; we applied EPLL to each channel using the kernel estimated for that channel. For the method in [5], we used the Matlab implementation provided at www.cs.nyu.edu/~dilip/wordpress/?page_id=159, again with kernel size of $(101,101)$.
@@ 5.3.2 Reduced Test Set
Due to the computational complexity of the EPLL deconvolution [26], as well as the deconvolution methods in Levin et al. [11] and Krishnan et al. [5], we generated reduced test datasets for these experiments. From our test dataset (see Section 3.3), we created two subsets to span length and angle. The first subset, subsequently referred to as Length 5 (L5), uses $5$ randomly selected blurred images for each length, totaling $500$ images. The second subset, subsequently referred to as Angle 3 (A3) uses $3$ randomly selected blurred images for each angle, totalling $540$ images. All images in these reduced test sets are noise free.

### v3

@@ 3.3 COCO Blurred Dataset
We use the 2014 COCO dataset [27] to create a blurred dataset for training and validating a model for blind estimation of length and angle given only a blurry image. The 2014 COCO dataset consists of a training dataset with 82,783 images and a validation set of 40,504 images. The training dataset is used to create a blurred training dataset and the validation dataset is split (with no overlap) to create blurred validation and test datasets.
Figure 4: Representation of COCO images versus length and angle. (a) Number of unique COCO images per angle in training and validation/testing datasets. (b) Number of unique COCO images per length in training and validation/testing datasets.
@@
The blurred dataset is created by defining a blur generator that loops through the labels $(r,\phi)_{\text{u}}$, creating the corresponding blur kernel using the skimage.draw function (see Sec. 3.2), and convolving the blur kernel with an image from the COCO dataset. A random COCO image is chosen (without replacement) for each length $r$ to provide a variety of images in the blurred dataset. There is a dataset imbalance in lengths and/or angles represented in the blurred dataset due to the uneven representation of angles for shorter lengths (see Fig. 3). We improve the balance by creating a minimum of 175 blurred images per length, implying that more unique COCO images are used for shorter blur lengths. Each run of the blur generator creates 21,789 blurred images. The creation of the training set used 12 parallel threads of the blur generator, each operating on a subset of the COCO training dataset, creating a total of 261,468 blurred images. The validation and testing sets were each created with one run of the blur generator operating on the validation and testing splits of COCO validation dataset, creating a total of 21,789 blurred images for each. Figure 4 shows the number of unique COCO images versus length and angle for the training, validation, and testing datasets. We notice a spike in the number of unique images for length $r=1$ and for angle $\phi=0$ due to the non-blurred images that are part of the dataset. In general, more unique images are used for shorter lengths as the blur generator needs to loop through more images to complete the minimum threshold of images per length. We notice also that small length blur kernels create peaks at angles of $\phi=\{-45,0,45,90\}$ since smaller length kernels are limited in the angles they are capable of representing.
Figure 5: Distribution of blurred images for angle and length. (a) Number of blurred images per angle. (b) Number of blurred images per length.
@@
There is an interdependence between angle and length, resulting in an inability to completely balance both length and angle in the blurred dataset. Figure 5 shows the distribution of labeled blurred images with respect to angle and length. Our choice to generate a minimum of 175 blurred images per length in the blur generator favors a more uniform length distribution [Fig. 5(b)]. This, however, results in peaks in the angle distribution [Fig. 5(a)] due to the over-representation of certain angles for smaller lengths (see Fig. 3). A choice to favor a more uniform angle distribution, however, would result in smaller representation of shorter blur lengths, creating a more problematic data imbalance. While there are other choices that could be made regarding data imbalance here, we find that this blurred dataset can be used to train a network to accurately predict both length and angle.
@@ 4 Uniform Blur Prediction
We formulate blur length and angle estimation as a deep learning regression problem using the VGG16 [6] architecture as the backbone of our model. VGG16 was trained from scratch with TensorFlow’s default layer initializations to predict both the length and angle parameters for a uniform linear motion blurred image. The last layer of VGG16 was modified to have two output nodes (corresponding to prediction of length and angle) with sigmoid activations. Since the sigmoid activation outputs values in the range $[0,1]$, length and angle are normalized to be in the range $[0,1]$ to train the model. The predicted length and angle parameters are re-scaled to their native ranges ($r\in[1,100]$, $\phi\in(-90,90]$) for validation.
Due to the fixed input size of VGG16 ($224\times 224$ pixels), we created a TensorFlow data generator to randomly crop the blurred COCO images. We crop instead of resizing the image to maintain accuracy in the blur angle labels since a resize could change the aspect ratio of the image and thus the blur angle. If a blurred image was smaller than the input size for VGG16, it was skipped and not used in training; in total there were 892 training, 5 validation, and 13 test images that were skipped. This results in less than $0.3\%$ of images skipped for each dataset.
We use the Adam optimizer to with a learning rate of 0.1 and epsilon of 0.1 and the mean squared error (MSE) as the loss function. A batch size of 50 was empirically determined to mitigate convergence issues. The model is trained for 50 epochs, saving the weights for the best model throughout training; training is terminated if the MSE performance has not improved within the previous 5 epochs. Training takes about 25 minutes per epoch and about 12 hours to fully train on an NVIDIA RTX-3090.
We performed multiple trainings of the network for different noise levels, resulting in four different models. The first model is trained with no noise and the other three models are trained with different levels of additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$, corresponding to signal-to-noise (SNR) values of $\{30,20,10\}$ dB, respectively. Noise is added in our data generator after the blurred image is cropped and normalized to the range $[0,1]$. Each model is tested by predicting blur parameters for the noiseless test set and for three additional test runs in which three noise levels are added to the test set.
@@ 5 Experiments and Results
This section presents results using the $R^{2}$ coefficient of determination metric for evaluating the regression model. Next, we consider a scenario with additive noise and present results our model’s predictions. Finally we compare with other methods of blur kernel estimation [9, 14, 25] and evaluate by assessing deblurred images using the error ratio score as introduced by [7].
@@ 5.1.1 Accuracy of Parameter Estimation
In testing, we measure performance using the coefficient of determination, $R^{2}$, which measures the goodness of fit between actual known values of variable $y_{i}$ and estimated values $x_{i}$:
@@
where $\bar{y}$ is the mean of the known variables $y_{i}$ and $n$ is the number of samples [29]. The numerator $\sum_{i=1}^{n}(y_{i}-x_{i})^{2}$ is the sum of squares of the residual prediction errors and the denominator $\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}$ is proportional to the variance of the known data. A perfect model will have zero residual errors and thus an $R^{2}=1$. A naïve model that always predicts the average of the data $\bar{y}$ will have equal numerator and denominator and thus an $R^{2}=0$. Models that have predictions worse than the naïve model will have an $R^{2}<0$.
@@ 5.1.2 Quality of Deblurred Image
[…]
We use the sum of squared differences (SSD) as in [7] to define the error ratio. Error ratios are generally presented as a cumulative histogram for error ratios binned in the range $[1,4]$. As noted in [7], SSD error ratios above 2 tend to indicate significant perceptually noticeable distortion present in the image deconvolved with the estimated kernel as compared to the image deconvolved with the true kernel.
@@ 5.2.1 Noise-Free Predictions
A VGG16 based model (Sec. 4) was first trained on blurred images without any additive noise. Scatter plots of estimated versus actual length and angle are shown in Fig. 6 along with the corresponding $R^{2}$ scores. We find the model to be highly accurate for prediction of both length and angle as demonstrated by the $R^{2}$ scores of 0.9869 and 0.9935, respectively. In Fig. 6(a) we note a larger spread in estimated values as the length increases. This is not surprising, as larger blur lengths are expected to be more difficult to accurately estimate. In Fig. 6(b) we note certain angles have a larger spread in estimated values. This is most notable for $\phi\in\{-45,0,45,90\}$, but can be noted for other angles. This is due to errors in prediction for the smaller length kernels which have a limited set of angles. It is important to recall, however, that many of these incorrect angle predictions will result in a correct blur kernel. As an example, an image blurred with kernel parameters $(r,\phi)=(2,0)$ may have an angle prediction of $\hat{\phi}=27$. However, since blur kernels of length $r=2$ can only be represented by $\phi\in\{-45,0,45,90\}$, the blur kernel created with parameters $(r,\phi)=(2,27)$ will generate a blur kernel estimate of $(r,\phi)_{\text{u}}=(2,0)$ which is correct.
Figure 6: Scatterplots of estimated versus actual values for (a) length and (b) angle. Individual estimates are represented as blue points and the best fit linear line is plotted in red. (a) Scatterplot of estimated versus actual length, $R^{2}=0.9869$. (b) Scatterplot of estimated versus actual angle, $R^{2}=0.9935$.
@@ 5.2.2 Predictions with Additive Noise
We used the model from Sec. 5.2.1, trained on noise-free blurred images, and tested it on images with additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$, corresponding to SNR values of $\{30,20,10\}$ dB, respectively. Results for this experiment are shown in the first row of Tables 1 and 2 for length and angle prediction, respectively. We note the estimation of the length parameter is more susceptible to noise than angle, resulting in a complete failure of prediction ($R^{2}$ scores less than 0) for even the smallest level of additive noise, $\sigma^{2}=0.001$. We hypothesize that additive noise can alter the intensity distribution along the blur path in a blurry image, creating the appearance of artificially shorter or longer blur paths. Those blur paths, however, will likely retain more characteristics of their angle for the same level of noise.
@@
We trained three additional models using noisy blurred images with additive white Gaussian noise with variance $\sigma^{2}\in\{0.001,0.01,0.1\}$ and tested each of those models on noise-free and noisy images, i.e., for $\sigma^{2}\in\{0,0.001,0.01,0.1\}$. Results for those three models are in the second through fourth rows of Tables 1 and 2 for length and angle prediction, respectively. We again note a higher sensitivity to noise in the prediction of length (Table 1) than angle (Table 2). We further note that the $R^{2}$ score decreases by $\sim 0.1$ for the model trained on the highest level of noise $\sigma^{2}=0.1$ and tested on noise-free images, compared to the noise-free model tested on noise-free images, but that the same model is robust to varying levels of noise. Finally we note that models trained on noisy data appear to be robust to noise levels less than or equal to the noise level on which they are trained. This implies that a single model trained on a single noise level can yield accurate predictions even for smaller noise levels not seen in training.
Most other methods [14, 11, 23, 17] that test under additive noise test up to a Gaussian noise of $\sigma^{2}=0.01$ to simulate sensor noise, while some [9, 18] test noise up to $\sigma^{2}=0.02$. In comparison, our model is tested up to $\sigma^{2}=0.1$ and demonstrates a higher tolerance for noise which can be an advantage when modeling atmospheric turbulence.
@@ 5.3.1 Deconvolution and Comparison Methods
We quantify the performance of our blur parameter estimation by using the expected patch log likelihood (EPLL) method [30] to deconvolve the images using our predicted motion blur kernel and the ground truth blur kernel. We used the Python implementation available at [31], replacing the Gaussian kernel with a linear motion blur kernel. EPLL only accepts square, odd-sized blur kernels, necessitating a padding of kernels to be square and odd-sized. We symmetrically zero pad the shortest side of the kernel to keep the kernel centered and pad to match the size of the longest side. If the longest side is even, however, this results in an even-sized square kernel. We use a similar approach to [32], creating four odd-sized kernels each zero padded with the line asymmetrically offset toward a different corner. The blurred image is deconvolved with each of the four kernels and the average of the four deconvolved images is considered the resulting deblurred image.
Additionally, we compare performance with the blur kernel estimation methods from Levin et al. [14], Krishnan et al. [9], and Carbajal et al. [25] These works also include methods for estimating the deblurred image. For fairness of comparisons, we use only the blur kernel estimate from [14, 9, 25] and use those blur kernel estimates in the same EPLL [30] deblurring framework as described above. For the method in [14], we used the Matlab implementation provided at [33], using the deconv_diagfe_filt_sps function with a kernel prediction size of $(101,101)$ to allow prediction of the largest kernel size in our blurred dataset. Since the method in [14] estimates a blur kernel for each of the three channels in a color image, we applied EPLL to each channel using the kernel estimated for that channel. For the method in [9], we used the Matlab implementation provided at [34], again with kernel size of $(101,101)$. For the method in [25], we used the python implementation provided at [35] with the fixed kernel size of $(33,33)$; we note that this will put this method at a disadvantage in predicting longer blur lengths. Additionally, we estimated a single blur kernel for the uniformly blurred images by using the image-averaged mixing coefficients in the superposition of the kernel bases; this allows us to use the same EPLL deblurring framework for comparison to the other methods.
@@ 5.3.2 Reduced Test Set
Due to the computational complexity of the EPLL deconvolution [30], as well as the kernel estimation methods in [9, 14, 25], we generated reduced-size test datasets for these experiments. From the test dataset (Sec. 3.3), two subsets are created to span length and angle, respectively. The first subset, subsequently referred to as Length 5 (L5), uses $5$ randomly selected blurred images for each length, totaling $500$ images. The second subset, subsequently referred to as Angle 3 (A3) uses $3$ randomly selected blurred images for each angle, totaling $540$ images. All images in these reduced test sets are noise free.